OpenStudy (kj4uts):

If h(x)=(fog)(x) and h(x)=5(x+1)^3, find f(x) and g(x) a. f(x)=5x^3, g(x)=(x+1)^3 b. f(x)=(x+1)^3, g(x)=5x^3 c. f(x)=x+1, g(x)=5x^3 d. f(x)=5x^3, g(x)=x+1 Please help and explain!

OpenStudy (math&ing001):

h(x) = (fog)(x) = f(g(x)) Say g(x) = x+1 then h(x) = f(g(x)) = 5(g(x))^3 So f(x) = 5x^3.

OpenStudy (math&ing001):

@KJ4UTS